OpenStudy (anonymous):

The equation of line CD is y = 3x − 3. Write an equation of a line perpendicular to line CD in slope-intercept form that contains point (3, 1).

OpenStudy (anonymous):

y = 3x + 0 y = −3x − 8 y = negative 1 over 3x + 2 y = 1 over 3x + 0

OpenStudy (anonymous):

if two lines are perpendicular then their slopes, m1 and m2, m1 = -1/m2

OpenStudy (anonymous):

so C? @bloopman

OpenStudy (anonymous):

mhm

OpenStudy (antonio_xx2):

do u mine if i tag some 1 in beautiful

OpenStudy (anonymous):

go ahead

OpenStudy (antonio_xx2):

ok;)

OpenStudy (antonio_xx2):

@dumbcow

OpenStudy (dumbcow):

looks like they got it @EmmaDrew94 , do you understand why answer is C

OpenStudy (anonymous):

yes @dumbcow